Uranium disintegrates at a rate3 proportional to the amount present at any instant. If m1 and m2 grams of uranium are present at time t1 and t2 , respectively, show that half life of uranium is (t1-t2)In2/In(m1/m2)
1
Expert's answer
2020-11-26T20:06:08-0500

"\nu=\frac mM=\frac{N}{N_A}," "m \sim M."

"\frac{m}{m_0}=2^{-\frac tT},"

"log_2=-\frac tT,"

"T=-\frac{t}{log_2\frac{m}{m_0}}=\frac{t}{log_2\frac{m_0}{m}}=\frac{tln2}{ln\frac{m_0}{m}},"

"\frac{lnm_0-lnm}{tln2}=\frac 1T,"

"lnm_0=\frac{tln2}{T}+lnm=const,"

"\frac{t_1ln2}{T}+lnm_1=\frac{t_2ln2}{T}+lnm_2,"

"\frac{(t_1-t_2)ln2}{T}=lnm_2-lnm_1,"

"T=\frac{(t_1-t_2)ln2}{ln\frac{m_2}{m_1}}."